The total sales for restaurant b was $450,000 more than restaurant b and the total sales were 550,000,000 what were the sales for the years


1
Expert's answer
2021-10-06T17:01:39-0400

"a-b=450000"

"a+b=550000000"

"a=\frac{550000000+450000}{2}=275,225,000"

"b=a-450000=275000000-450000=274,775,000"